### v3

# Similarity of Neural Architectures using Adversarial Attack Transferability

Abstract
In recent years, many deep neural architectures have been developed for image classification. Whether they are similar or dissimilar and what factors contribute to their (dis)similarities remains curious. To address this question, we aim to design a quantitative and scalable similarity measure between neural architectures. We propose Similarity by Attack Transferability (SAT) from the observation that adversarial attack transferability contains information related to input gradients and decision boundaries widely used to understand model behaviors. We conduct a large-scale analysis on 69 state-of-the-art ImageNet classifiers using our proposed similarity function to answer the question. Moreover, we observe neural architecture-related phenomena using model similarity that model diversity can lead to better performance on model ensembles and knowledge distillation under specific conditions. Our results provide insights into why developing diverse neural architectures with distinct components is necessary.
1 Introduction
The advances in the deep neural networks (DNN) architecture design have taken a key role in their success by making the learning process easier (e.g., normalization [50, 109, 2] or skip connection [40]), enforcing human inductive bias in the architecture [57], or increasing model capability with the self-attention mechanism [101]. With different architectural components containing architectural design principles and elements, a number of different neural architectures have been proposed. They have different accuracies, but several researches have pointed out that their predictions are not significantly different [69, 68, 33].
By this, can we say that recently developed DNN models with different architectural components are similar or the same? The answer is no. It is because a model prediction is not the only characteristic to compare their similarities. Existing studies have found differences by focusing on different features, such as layer-by-layer network component [55, 77], a high-level understanding by visualization of loss surface [27], input gradient [86, 84], and decision boundary [85]. Researchers could understand the similarity between models through these trials; however, the similarity comparison methods from previous studies are insufficient for facilitating comprehensive studies because they do not satisfy two criteria that practical metrics should meet: (1) providing a quantitative similarity score and (2) being compatible with different base architectures (e.g., CNN and Transformer). Recently, Somepalli et al. [85] suggested a quantitative similarity metric based on the on-manifold decision boundary. However, this method needs heavy computation due to the triplet sampling process when the number of classes becomes large (e.g., ImageNet dataset with 1,000 classes). If we limit the total computations, this method will suffer from inaccurate approximation with high variance.
Figure 1: t-SNE plot showing 10 clusters of 69 neural networks using our similarity function, SAT. We cluster 69 neural networks in 10 groups based on SAT. The group number follows the result of spectral clustering (Tab. 2). Due to the space limit, we only illustrate representative models in each group.
This paper proposes a quantitative similarity function that is scalable and easily applicable to diverse architectures, named Similarity by Attack Transferability (SAT). We focus on adversarial attack transferability (AT), which indicates how generated adversarial perturbation is transferable between two different architectures. Specifically, it is widely studied that the vulnerable points of neural architectures depend on their own architectural property or how models capture the features from inputs, such as the usage of self-attention [31], the stem layer [48], and the model dependency on high or low-frequency components of input [3, 54]. Thus, if two different neural architectures are similar, the AT between the models is high because they share similar vulnerability [79]. Furthermore, AT can be a reliable approximation for comparing the input gradients [67], decision boundary [53], and loss landscape [25] of models. All of them are widely-used frameworks to understand model behavior and the difference between models and used to measure the similarity of models in previous works [90, 83, 86, 84, 89, 4, 16, 85, 27, 61, 73]; in other words, we can capture various model characteristics by SAT.
We quantitatively measure pairwise SATs of 69 different ImageNet-trained neural architectures, provided by Wightman [107]. With SAT similarity scores of 69 architectures, we analyze what components among 13 architectural components (e.g., normalization, activation, the design choice for stem layers) that consist of neural architectures largely affect model diversity by gradient boosting regression and clustering on model similarities. Furthermore, we observe relationships between model similarity and model ensemble performances, and distillation performances. The main contributions of this work can be summarized as follows.
  • We propose a scalable and quantitative similarity between neural architectures, named Similarity by Attack Transferability (SAT). SAT is easy to conduct and reports stable results even when there is a change in data sampling.
  • We conduct a comprehensive analysis using 69 models to investigate what components make architectures behave similarly or differently. To the best of our knowledge, this is the first large-scale analysis exploring the relationship between model similarity and architectural components. Furthermore, our findings are not limited to architectural components but also encompass training strategies; different training strategies do not affect model diversity as much as architectural differences.
  • We identify several outgrowths using SAT. First, we find model dissimilarity leads to better ensemble performance and analyze the reasons behind this. Additionally, we suggest new guidelines for selecting a teacher model suited to the student model in knowledge distillation. These observations provide insights into why we need to develop diverse architectures.
2 Related work
Similarity between DNNs has been actively explored recently. Several studies focused on comparing intermediate features for understanding the behavior of DNNs. Raghu et al. [77] observed the difference between layers, training methods, and architectures (e.g., CNN and ViT) based on layer-by-layer comparison [55]. Some studies have focused on loss landscapes, which helps to visualize the loss of models on the parameter space [27, 61, 73]. Although these methods show a visual inspection, they cannot support quantitative measurements. In this paper, we aim to support a quantitative similarity function.
Another line of research has been focused on prediction-based statistics, e.g., comparing wrong and correct predictions [58, 32, 33, 81]. However, as recent complex DNNs are getting almost perfect, just focusing on prediction values can be misleading. For example, Meding et al. [69] observed that recent DNNs show highly similar predictions. In this case, prediction-based methods will be no more informative. On the other hand, our SAT can provide meaningful information and findings for 69 recent neural architectures.
Input gradient is another popular framework to understand model behavior by observing how a model will change predictions by local pixel changes [83, 86, 84, 89, 4]. If two models are similar, their input gradients will also be similar. These methods are computationally efficient, and no additional training is required; they can provide a visual understanding of the given input. However, input gradients are inherently noisy; thus, these methods will need additional pre-processing, such as smoothing, for a stable computation [16]. Also, input gradient-based quantity measures usually measure how the input gradient matches the actual foreground, i.e., we need ground-truth foreground masks for measuring such scores. On the contrary, SAT needs no additional pre-processing and mask annotations.
Comparing the decision boundaries of models will provide a high-level understanding of how models behave differently for input changes and how models extract features from complicated data dimensions. Somepalli et al. [85] analyzed characteristics of models by comparing their decision boundaries on the on-manifold plane constructed by three random images. However, we observe that this approach derives unstable measurements on huge datasets (e.g., ImageNet) and is expensive compared to SAT.
3 Similarity by Attack Transferability (SAT)
In this section, we propose a quantitative similarity measure between two neural architectures using adversarial attack transferability, which indicates whether an adversarial sample from a model can fool another model or not. The concept of adversarial attack has effectively pointed out the vulnerabilities of DNNs by input gradient [91, 34, 67].
Interestingly, these vulnerabilities have been observed to be intricately linked to architectural properties. Fu et al. [31] demonstrated the effect of the attention modules in architecture in terms of attack success rate. Hwang et al. [48] analyzed that the stem layer structure of models causes models to have different adversarial vulnerable points in the input space, e.g., video action-recognition models periodically have vulnerable frames, such as every four frames. In other words, an adversarial sample to a model highly depends on the inherent architectural property of the model, e.g., the base architecture or the stem layer structure.
Another perspective emphasized the dissimilarities in dependencies on high-frequency and low-frequency components between CNN-based and transformer-based models, showing different vulnerabilities to different adversarial attacks [3, 54]. As shown by Park and Kim [73], different architectural choices behave as different frequency filters (e.g., the self-attention mechanism works as a low-pass filter, while the convolution operation works as a high-pass filter); thus, we can expect that the different architectural component choice will affect to the model vulnerability, e.g., vulnerability to high-frequency perturbations. In other words, if we can measure how the adversarial vulnerabilities of the models are different, we also can measure how the networks are similar or dissimilar.
To measure how model vulnerabilities differ, we employ adversarial transferability, where it indicates whether an adversarial sample from a model can fool another model. If two models are more similar, their adversarial transferability gets higher [79, 25, 62]. On the other hand, because the adversarial attack targets vulnerable points varying by architectural components of DNNs [47, 31, 48, 54], if two different models are dissimilar, the adversarial transferability between them gets lower. Furthermore, attack transferability can be a good approximation for measuring the differences in input gradients [67], decision boundaries [53], and loss landscape [25], where they are widely used techniques for understanding model behavior and similarity between models as discussed in the related work section. While previous approaches are limited to non-quantitative analysis, inherent noisy property, and computational costs, adversarial transferability can provide quantitative measures with low variances and low computational costs.
We propose a new similarity function that utilizes attack transferability, named Similarity by Attack Transferability (SAT), providing a reliable, easy-to-conduct, and scalable method for measuring the similarity between neural architectures. Formally, we generate adversarial samples $x_{A}$ and $x_{B}$ of model $A$ and $B$ for the given input $x$. Then, we measure the accuracy of model $A$ using the adversarial sample for model $B$ (called $\text{acc}_{B\rightarrow A}$). If $A$ and $B$ are the same, then $\text{acc}_{B\rightarrow A}$ will be zero if the adversary can fool model B perfectly. On the other hand, if the input gradients of $A$ and $B$ differ significantly, then the performance drop will be neglectable because the adversarial sample is almost similar to the original image (i.e., $\|x-x_{B}\|\leq\varepsilon$). Let $X_{AB}$ be the set of inputs where both $A$ and $B$ predict correctly, $y$ be the ground truth label, and $\mathbb{I}(\cdot)$ be the indicator function. We measure SAT between two different models by:
$\text{SAT}(A,B)=\log\big{[}\max\big{\{}\varepsilon_{s},100\times$ (1)
$\frac{1}{2|X_{AB}|}\sum_{x\in X_{AB}}\left\{\mathbb{I}({A(x_{B})}\neq y)+\mathbb{I}({B(x_{A})}\neq y)\right\}\big{\}}\big{]},$
where $\varepsilon_{s}$ is a small scalar value to prevent the ill-posedness of $\text{SAT}(\cdot)$. If $A=B$ and we have an oracle adversary, then $\text{SAT}(A,A)=\log 100$. In practice, a strong adversary (e.g., PGD [67] or AutoAttack [22]) can easily achieve a nearly-zero accuracy if a model is not trained by an adversarial attack-aware strategy [67, 21]. On the other hand, if the adversarial attacks on $A$ are not transferable to $B$ and vice versa, then $\text{SAT}(A,B)=\log\varepsilon_{s}$.
SAT (Eq. 1) can act as a similarity function between models. Ideally, we would like to define a similarity function $d$ between two models with the following three properties: (1) $n=\arg\min_{m}d(n,m)$, (2) $d(n,m)=d(m,n)$ and (3) $d(n,m)>d(n,n)$ if $n\neq m$. The proposed function satisfies these properties if we assume the perfect adversary. If the adversary is perfect, then $\text{acc}_{A\rightarrow A}$ will be zero, and it will be the minimum because accuracy is non-negative. “$\text{acc}_{A\rightarrow B}+\text{acc}_{B\rightarrow A}$” is symmetric thereby SAT is symmetric. Finally, SAT will satisfy $d(n,m)\geq d(n,n)$ if $n\neq m$ where it is a weaker condition than (3).
Now, we extensively analyze 69 recent state-of-the-art ImageNet-trained networks using our SAT score by focusing on two questions. (1) Which network component contributes to the diversity between models? (2) Why do we need to develop various neural architectures?
Settings. We use the PGD attack for the adversary. We set the iteration to 50, the learning rate to 0.1, and $\varepsilon$ to $8/255$. We select 69 neural architectures trained on ImageNet [80] from the PyTorch Image Models library [107]. To reduce the unexpected effect of a significant accuracy gap, the chosen model candidates are limited to the models whose top-1 accuracy is between 79% and 83%. We also ignore the models with unusual training techniques, such as training on extra training datasets, using a small or large input resolution (e.g., less than 200 or larger than 300), or knowledge distillation. When $A$ and $B$ take different input resolutions, then we resize the attacked image from the source network for the target network. We also sub-sample 10% ImageNet validation images (i.e., 5,000 images) to measure the similarity. This strategy makes our similarity score more computationally efficient. The full pairwise similarities of 69 models can be found in Appendix.
Ablation study on attack methods. In our analyses, we compute SAT using the famous PGD attack. To explore the effect of the choice of attack methods, we compute SAT using two different gradient-based attack methods; Autoattack [22], the state-of-the-art attack method based on the ensemble of four diverse attacks, and Patchfool attack [31], specially designed for vision transformers. We sample 8 representative models among 69 models for testing the effect of Autoattack on SAT: ViT-S, CoaT-Lite Small, ResNet-101, LamHaloBotNet50, ReXNet-150, NFNet-L0, Swin-T and Twins-pcpvt. Fig. 1(a) shows the high correlation between SAT scores using PGD and Autoattack; it shows a correlation coefficient of 0.98 with a p-value of $1.43\times 10^{-18}$. For testing the Patchfool attack, we only generate adversarial perturbations on ViT-S and get attack transferability to all other models (68 models) because it only targets Transformers. In Fig. 1(b), Patchfool also shows a high correlation compared to the PGD attack (correlation coefficient 0.91 with p-value $3.62\times 10^{-27}$). The results show that SAT score is robust to the choice of attack methods if the attack is strong enough. Note that PatchFool needs a heavy modification on the model code to extract attention layer outputs manually. Contrarily, PGD and AutoAttack are model-agnostic and need no model code modification. Thus, as PatchFool and PGD show similar rankings, we use PGD due to its simplicity.
(a) PGD vs. Autoattack
(b) PGD vs. Patchfool
Figure 2: Effect of different adversarial attack methods to SAT. The trend line and its 90% confidence interval are shown. We show the relationship between our SAT using PGD and SAT using other attacks, (a) Autoattack and (b) Patchfool attack.
Comparison with other similarities. Somepalli et al. [85] proposed a sampling-based similarity score for comparing decision boundaries of models. SAT has two advantages over Somepalli et al.: computational efficiency and the reliability of the results. First, SAT involves sampling 5,000 images and using 50-step PGD; the computation cost is [5,000 (sampled images) $\times$ 50 (PGD steps) + 5,000 (test to the other model)] $\times$ 2 (two models). On the other hand, Somepalli et al. [85] sample 500 triplets and generate 2,500 points to construct decision boundaries. In this case, the whole amount of inference is [500 (sampled triplets) $\times$ 2,500 (grid points)] $\times$ 2 (two models), which is 4.9 times larger than our method. Secondly, Somepalli et al. [85] sampled three images of different classes. As the original paper used the CIFAR-10 dataset [56], 500 triplets can cover all possible combinations of three classes among the ten classes (($10\choose 3$ = 120 $<$ 500). However, in datasets with many classes, it becomes computationally infeasible to represent all possible combinations of three classes (e.g., ImageNet needs ($1000\choose 3$ = 164,335,500, almost 130 times greater than the number of training examples of ImageNet). Additionally, we find that the similarity results are unreliable (i.e., showing a high variance) when the number of sample triplets is small. In Tab. 1, we calculate the similarity scores between ConvNeXt-T [66] and Swin-T [65] from ten different sets with varying sample sizes. SAT exhibits significantly better stability (i.e., low variances) than Somepalli et al. Note that we use our similarity measurement as the percentage degree without the logarithmic function and control the scale of samples to maintain similar computation complexity between SAT and the compared method for a fair comparison.
Somepalli et al. [85] SAT (ours)
# triplets std cost # images std cost
10 (4.8×$\times 10^{-13}$) 4.49 1.0 500 (0.01) 1.88 1.0
20 (9.6×$\times 10^{-13}$) 3.28 2.0 1000 (0.02) 1.05 2.0
50 (2.4×$\times 10^{-12}$) 1.63 5.0 2500 (0.05) 0.91 5.2
100 (4.8×$\times 10^{-12}$) 1.54 10.0 5000 (0.1) 0.77 10.2
Table 1: Comparison of stability of measurements. We compare the stability of method by [85] and our method. Stability is indicated by the standard deviation (std). The numbers in ($\cdot$) mean the sampling ratio to all possible combinations to compute the exact value. “cost” denotes relative costs compared to the total forward costs for the 50K ImageNet validation set: Somepalli et al. needs ($50\text{K}\choose 3$ = 2.1×$\times 10^{13}$ and SAT needs 50K. Here, we assume that forward and backward computations cost the same.
4 Model analysis by network similarity
We quantitatively analyze how each model component affects the similarity or diversity between networks by feature importance analysis and clustering analysis. We also show that the diversity caused by different training strategies is not as significant as caused by different neural architectures.
4.1 Which architectural component causes the difference between architectures?
Figure 3: Importance of architectural components to network similarity. 13 components are sorted by the contribution to the network similarities. The larger feature importance means the component contributes more to the network similarity.
[…]
Feature important analysis. The full list of components of 69 architectures is in Appendix. Now, we measure the feature importance by fitting a gradient boosting regressor [30] on the feature difference (e.g., $f_{\text{ResNet-50}}-f_{\text{DeiT-base}}$) measured by Hamming distance and the corresponding similarity. The details of the regressor are described in Appendix. We use the permutation importance [7] that indicates how the trained regression model changes the prediction according to randomly changing each feature. The feature importance of each architectural component is shown in Fig. 3. We first observe that the choice of base architecture (e.g., CNN [57], Transformer [101], and MLP-Mixer [95]) contributes to the similarity most significantly. Fig. 3 also shows that the design choice of the input layer (i.e., stem layer design choice or input resolution) affects the similarity as much as the choice of basic components such as normalization layers, activation functions, and the existence of attention layers. On the other hand, we observe that the modified efficiency-aware convolution operations, such as depth-wise convolution [17] or group convolution [110], are ineffective for model diversity.
No. Top-5 Keywords Architecture
1 Stem layer: 16$\times$16 conv w/ s16, ConViT-B [29], CrossViT-B [11],DeiT-B [96], DeiT-S [96],
No Hierarchical, GeLU, LN, Final GAP ViT-S (patch size 16) [28],ResMLP-S24 [97], gMLP-S [64]
2 Stem layer: 4$\times$4 conv w/ s4, LN, GeLU, Twins-PCPVT-B [18], Twins-SVT-S [18], CoaT-Lite Small [24],
Transformer, No pooling at stem NesT-T [121], Swin-T [65], S3 (Swin-T) [12], ConvNeXt-T [66], ResMLP-B24 [97]
3 Transformer, Final GAP, GeLU, XCiT-M24 [1], XCiT-T12 [1], HaloRegNetZ-B [^2] Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., TNT-S [39],
Pooling at stem, InRes: 224 Visformer-S [15], PiT-S [44], PiT-B [44]
4 Stem layer: stack of 3$\times$3 conv, 2D SA, HaloNet-50 [102], LambdaResNet-50 [5], BoTNeT-26 [87],
InRes: 256, Pooling at stem, SiLU GC-ResNeXt-50 [10], ECAHaloNeXt-50222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., ECA-BoTNeXt-26222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63].
5 Stem layer: stack of 3$\times$3 convs, InRes: 256, LamHaloBoTNet-50222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., SE-BoTNet-33222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., SE-HaloNet-33222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63].,
2D SA, CWA: middle of blocks, CNN Halo2BoTNet-50222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., GC-ResNet-50 [10], ECA-Net-33 [106]
6 Stem layer: 7$\times$7 conv w/ s2, ReLU, ResNet-50 [40], ResNet-101 [40], ResNeXt-50 [110], Wide ResNet-50 [116],
Pooling at stem, CNN, BN SE-ResNet-50 [46], SE-ResNeXt-50 [46], ResNet-V2-50 [41], ResNet-V2-101 [41],
ResNet-50 (GN) [109], ResNet-50 (BlurPool) [120], DPN-107 [14], Xception-65 [17]
7 NAS, Stem layer: 3$\times$3 conv w/ s2 EfficientNet-B2 [93], FBNetV3-G [23], ReXNet ($\times$1.5) [38], RegNetY-32 [76],
CWA: middle of blocks, CWA, DW Conv MixNet-XL [94], NF-RegNet-B1 [8]
8 Input resolution: 224, Stem layer: stack of 3$\times$3 convs, NFNet-L0222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., ECA-NFNet-L0222Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63]., PoolFormer-M48 [113], ResNeSt-50 [118], ResNet-V2-50-D-EVOS [^2] Customized models by [107]: HaloRegNetZ = HaloNet + RegNetZ; ECA-BoTNeXt = ECA-Net + HaloNet + ResNeXt; ECA-BoTNeXt = ECA-Net + BoTNet + ResNeXt; LamHaloBoTNet = LambdaNet + HaloNet + BoTNet; SE-BoTNet = SENet + BoTNet; SE-HaloNet = SENet + HaloNet; Halo2BoTNet = HaloNet + BoTNet; NFNet-L0 = an efficient variant of NFNet-F0 [9]; ECA-NFNet-L0 = ECA-Net + NFNet-L0; ResNet-V2-D-EVOS = ResNet-V2 + EvoNorms [63].,
Group Conv, Final GAP, 2D SA ConvMixer-1536/20 [99]
9 ReLU, Input resolution: 224, DW Conv, BN, ViT-B (patch size 32) [28], R26+ViT-S [88], DLA-X-102 [112], eSE-VoVNet-39 [60],
2D self-attention ResNet-101-C [43], RegNetX-320 [76], HRNet-W32 [105]
10 ReLU + Leaky ReLU, InRes: 256, CSPResNet-50 [104], CSPResNeXt-50 [104], CSPDarkNet-53 [6], NF-ResNet-50 [8]
Stem layer: 7$\times$7 conv, CNN, Pooling at stem
Table 2: Clustering by similarities. We perform a clustering analysis on 69 architectures. All the architectures here are denoted by the aliases defined in their respective papers. We show the top-5 keywords for each cluster based on TF-IDF. InRes, SA, and CWA denote input resolution, self-attention, and channel-wise attention, respectively. The customized model details are described in the ${}^{\dagger}$.
Clustering analysis. We additionally provide a clustering analysis based on the architectural similarities. We construct a pairwise similarity graph with adjacency matrix $A$ between all 69 architectures where its vertex denotes an architecture, and its edge denotes the similarity between two networks. We perform the spectral clustering [71] on $A$ where the number of clusters $K$ is set to 10: We compute the Laplacian matrix of $A$, $L=D-A$ where $D$ is the diagonal matrix and its $i$-th component is $\sum_{j}A_{ij}$. Then, we perform K-means clustering on the $K$-largest eigenvectors of $L$, where $K$ is the number of clusters. The pairwise distances of spectral features (i.e., 10-largest eigenvectors of $L$) of 69 neural architectures are shown in Fig. 3(a). The rows and columns of Fig. 3(a) are sorted by the clustering index (Tab. 2). We can see the block-diagonal patterns, i.e., in-clusters similarities are more significant than between-clusters similarities. More details are in Appendix.
Tab. 2 shows the clustering results on 69 networks and the top-5 keywords for each cluster based on term frequency-inverse document frequency (TF-IDF) analysis. Specifically, we treat each model feature as a word and compute TF and IDF by treating each architecture as a document. Then we compute the average TF-IDF for each cluster and report top-5 keywords. Similar to Fig. 3, the base architecture (e.g., CNN in Cluster 5, 6, 10 and Transformer in Cluster 2, 3) and the design choice for the stem layer (e.g., Cluster 1, 2, 4, 5, 6, 7, 8, 10) repeatedly appear at the top keywords. Especially, we can observe that the differences in base architecture significantly cause the diversity in model similarities, e.g., non-hierarchical Transformers (Cluster 1), hierarchical networks with the patchification stem (Cluster 2), hierarchical Transformers (Cluster 3), CNNs with 2D self-attention (Cluster 4, 5), ResNet-based architectures (Cluster 6), and NAS-based architectures (Cluster 7).
(a) Spectral feature
(b) Similarity score
Figure 4: Spectral feature from SAT of ResNet-50 and ViT-s with four different optimization settings. The rows and columns of (a) are sorted by the clustering index. More details with model names are described in Appendix. Init, LR, and WD are randomly chosen from models trained with different settings of initialization, learning rate, and weight decay in Tab. 3.
4.2 Training strategy and similarity score
The architectural difference is not the only cause of the model diversity. We compare the impact by different architecture choices (e.g., ResNet and ViT) and by different training strategies while fixing the model architecture, as follows: Different initializations can affect the model training by the nature of the stochasticity of the training procedure. For example, Somepalli et al. [85] showed that the decision boundary of each architecture could vary by different initializations. We also consider different optimization hyper-parameters (e.g., learning rate, weight decay). Finally, we study the effect of different training regimes (e.g., augmentations, type of supervision). For example, the choice of data augmentation (e.g., Mixup [117], or CutMix [114]) or label smoothing [92] can theoretically or empirically affect adversarial robustness [119, 72, 82, 19]. Similarly, we investigate the effect of different types of supervision, such as self-supervision (e.g., MoCo v3 [13], MAE [42] and BYOL [35]) or semi-weakly supervised learning [111]. Note that the training strategies mentioned later inevitably contain the earlier-mentioned training strategies. For example, when we train models with different training regimes, models have different initialization seeds and different optimization hyper-parameters. Comparing 69 different architectures also contains the effect of different initialization and optimization hyper-parameters and parts of different training regimes. This is necessary for achieving high classification performance. Experimental details about the used models are provided in Appendix.
Architecture Diversity by SAT
ResNet-50 Initialization 4.23
Hyper-parameter 4.05
Training regime 3.27
ViT-S Initialization 4.21
Hyper-parameter 4.22
Training regime 3.44
All 2.73
Table 3: SAT within the same architecture. We compare the average similarity within the same architecture but trained with different procedures, e.g., random initializations, optimization hyper-parameters, and different training regimes. We report the similarities of two neural architectures, ResNet-50 and ViT-S. “All” denotes the average similarity of 69 neural architectures.
Tab. 3 shows the comparison of similarity scores between the same architecture but different learning methods (a smaller similarity means more diversity). First, we observe that using different random initialization or different optimization hyper-parameters shows high correlations with each other (almost $\geq$ 4.2) while the average similarity score between various neural architectures is 2.73. In other words, the difference in initializations or optimization hyper-parameters does not significantly contribute to the model diversity. Second, we observe that using different learning techniques (e.g., different augmentation methods and types of supervision) remarkably affects the similarity score (3.27 for ResNet and 3.44 for ViT), but is not as significant as the architectural difference (2.73). Furthermore, the change of similarity caused by different initializations or hyper-parameters is less marked than the change caused by different architecture (Fig. 3(b)). These observations provide two insights into model similarity. First, they demonstrate that the diversity resulting from various training strategies is not significant enough to interfere with observing the diversity between various architectures. Second, constructing new architecture is a more efficient way to achieve varied models rather than re-training the same one. In the next section, we will show the need for diversity. The need for diversity is observed in the next section.
5 Analysis for Ensemble and Distillation
In this section, we analyze how our neural architecture similarity is related to downstream tasks involving more than one model. First, we show that using more diverse models will lead to better ensemble performance. Second, we study the relationship between knowledge distillation and SAT. Furthermore, we can suggest a similarity-based guideline for choosing a teacher model when distilling to a specific architecture. Through these observations, we can provide insights into the necessity of diverse models.
5.1 Model diversity and ensemble performance
Figure 5: Correlation between SAT and ensemble performance. The trend line and its 90% confidence interval are shown.
[…]
We first measure the 2-ensemble performances among the 69 neural architectures (i.e., the number of ensembles is $\binom{69}{2}=2346$). We plot the relationship between the model similarity and ERR in Fig. 5. We observe that there exists a strong negative correlation between the model similarity and the ensemble performance (Pearson correlation coefficient $-0.31$ with p-value $1.74\times 10^{-54}$), i.e., more diversity leads to better ensemble performance.
We also conduct $N$-ensemble experiments with $N\geq 2$ based on the clustering results (Tab. 2). We evaluate the average ERR of the ensemble of models from $k$ clusters, i.e., if $N=5$ and $k=3$, the ensembled models are only sampled from the selected 3 clusters while ignoring the other 7 clusters. We investigate the effect of model diversity and ensemble performance by examining $k=1\ldots N$ (i.e., larger $k$ denotes more diverse ensembled models). We report the result with the metric, error reduction rate (higher is better) in Fig. 5(a). ImageNet top-1 error of the ensemble is also shown in Appendix. In all metrics, we observe that the ensemble of more diverse models (i.e., by increasing the number of clusters for the model selection) shows better ensemble performance. Interestingly, in Fig. 5(a), we observe that when the number of clusters for the model selection ($k$) is decreased, the ensemble performance by the number of ensembled models ($N$) quickly reaches saturation. Similarly, in Fig. 5(b), we observe that the number of wrong samples by all models is decreased by selecting more diverse models. In other words, the different architectural choices can lead to different model decision boundaries.
(a) Error reduction rate
(b) “All wrong” samples ratio
Figure 6: Model diversity by architecture difference and ensemble performance. We report ensemble performances by varying the number of ensembled models ($N$) and the diversity of the ensembled models. The diversity is controlled by choosing the models from $k$ different clusters (larger $k$ denotes more diversity).
(a) ResNet-50
(b) ViT-S
Figure 7: Model diversity by training techniques and ensemble performance. The same metrics to Fig. 6 for ResNet-50 and ViT-S models in Tab. 3 are shown. “All” means $N$=$k$ in Fig. 6 (i.e., # of ensembled models = # of Clusters).
(a) Results on various teachers
(b) Transformer teachers
(c) CNN & other teachers
Figure 8: Model diversity and distillation performance. (a) We show the relationship between teacher-student similarity and distillation performance of 25 DeiT-S models distilled by various teacher networks. We show the relationship when the teacher and student networks are based on (b) Transformer and (c) otherwise.
Tab. 3 showed that the different training strategies are not as effective as different architectures in terms of diversity. To examine the effect of this observation on the ensemble scenario, we report the ensemble results of different training strategies, i.e., the same ResNet-50 and ViT-S models in Tab. 3. For comparison with different architectures, we also report the ensemble of different architectures where all ensembled models are from different clusters (i.e., when $N=k$ in Fig. 6). Fig. 7 shows that although using diverse training regimes (blue lines) improves ensemble performance most among other training techniques (red and green lines), the improvements by using different architectures (yellow lines) are more significant than the improvements by using different training regimes (blue lines) with large gaps.
5.2 Model diversity and knowledge distillation
Knowledge distillation [45] is a training method for transferring rich knowledge of a well-trained teacher network. Intuitively, distillation performance affects a lot by choice of the teacher network; however, the relationship between similarity and distillation performance has not yet been explored enough, especially for ViT. This subsection investigates how the similarity between teacher and student networks contributes to the distillation performance. There are several studies showing two contradictory conclusions; Jin et al. [51] and Mirzadeh et al. [70] showed that a similar teacher leads to better distillation performance; Touvron et al. [96] reports that distillation from a substantially different teacher is beneficial for ViT training.
We train 25 ViT-Ti models with different teacher networks from 69 models that we used by the hard distillation strategy [45]. Experimental details are described in Appendix. Fig. 7(a) illustrates the relationship between the teacher-student similarity and the distillation performance. Fig. 7(a) tends to show a not significant negative correlation between teacher-student similarity and distillation performance (--0.32 Pearson correlation coefficient with 0.12 p-value). However, if we only focus on when the teacher and student networks are based on the same architecture (i.e., Transformer), we can observe a strong positive correlation (Fig. 7(b)) – 0.70 Pearson correlation coefficient with 0.078 p-value. In this case, our observation is aligned with [51, 70]: a teacher similar to the student improves distillation performance. However, when the teacher and student networks are based on different architectures (e.g., CNN), then we can observe a stronger negative correlation (Fig. 7(c)) with --0.51 Pearson correlation coefficient and 0.030 p-value. In this case, a more dissimilar teacher leads to better distillation performance. We also test other factors that can affect distillation performance in Appendix; We observe that distillation performance is not correlated to teacher accuracy in our experiments.
Why do we observe contradictory results for Transformer teachers (Fig. 7(b)) and other teachers (Fig. 7(c))? Here, we conjecture that when the teacher and student networks differ significantly, distillation works as a strong regularizer. In this case, using a more dissimilar teacher can be considered a stronger regularizer (Fig. 7(c)). On the other hand, we conjecture that if two networks are similar, then distillation works as easy-to-follow supervision for the student network. In this case, a more similar teacher will work better because a more similar teacher will provide more easy-to-follow supervision for the student network (Fig. 7(b)). Our experiments show that the regularization effect improves distillation performance better than easy-to-follow supervision (i.e., the best-performing distillation result is by a CNN teacher). Therefore, in practice, we recommend using a significantly different teacher network for achieving better distillation performance (e.g., using RegNet [76] teacher for ViT student as [96]).
6 Limitations and discussions
We include three discussions related to SAT. Due to the space limit, we leave more details in Appendix. Our discussions include: (1) How adversarially trained models affect SAT score. (2) An efficient approximation of our analysis for a novel model (i.e., for a novel 70-th model for our analysis). Instead of generating adversarial attacks for all 70 models, just using attacked images from the novel model could be enough. (3) Other possible applications of SAT, e.g., a novel architecture development, analyzing the hyper-parameter sensitivity, model selection for diverse models.
7 Conclusion
We have explored similarities between image classification models to investigate what makes the model similar or diverse and whether developing and using diverse models is required. For quantitative and model-agnostic similarity assessment, we have suggested a new similarity function, named SAT, based on attack transferability demonstrating differences in input gradients and decision boundaries. Using SAT, we conduct a large-scale and extensive analysis using 69 state-of-the-art ImageNet models. We have shown that macroscopic architectural properties, such as base architecture and stem architecture, have a more significant impact on similarity than microscopic operations, such as types of convolution, with numerical analysis. Furthermore, the effectiveness of the training strategy is minor compared to model architectural designs. Finally, we have first discovered the advantages of using diverse models in ensemble or distillation strategies for a practical scenario with a large-scale training dataset and a highly complex architecture.
Author Contributions
This work is done as an internship project by J Hwang under the supervision of S Chun. S Chun initialized the project idea: understanding how different architectures behave differently by using an adversarial attack. J Hwang, S Chun, and D Han jointly designed the analysis toolbox. J Hwang implemented the analysis toolbox and conducted the experiments with input from S Chun, D Han, and J Lee. J Hwang, D Han, B Heo, and S Chun contributed to interpreting and understanding various neural architectures under our toolbox. The initial version of “model card” (Tab. A.3 and A.4) was built by J Hwang, S Park, and verified by D Han and B Heo. B Heo contributed to interpreting distillation results. All ResNet and ViT models newly trained in this work were trained by S Park. J Lee supervised J Hwang and verified the main idea and experiments during the project. S Chun and J Hwang wrote the initial version of the manuscript. All authors contributed to the final manuscript.
Supplementary Material
We include additional materials in this document, such as the network details, feature importance analysis details, spectral clustering method, training settings, and setting of knowledge distillation. We also provide experimental results of spectral clustering, computing pairwise similarities, ensemble performance, and distillation performance. We discuss some limitations and discussion points of our work, such as the effect of an efficient approximation, and possible applications.
Appendix A Details of networks used for our analysis
We use 69 models in our research to evaluate the similarity between models and to investigate the impact of model diversity. In the main paper, we mark the names of models based on their research paper and PyTorch Image Models library (timm; 0.6.7 version) [107]. Tab. A.1 shows the full list of the models based on their research paper and timm alias.
In Tab. A.2, we show brief information about the architectural components. The full network specification is shown in Tab. A.3 and Tab. A.4. We follow the corresponding paper and timm library to list the model features.
in timm in paper in timm in paper in timm in paper
botnet26t_256 BoTNet-26 gluon_xception65 Xception-65 resnet50_gn ResNet-50 (GN)
coat_lite_small CoaT-Lite Small gmlp_s16_224 gMLP-S resnetblur50 ResNet-50 (BlurPool)
convit_base ConViT-B halo2botnet50ts_256 Halo2BoTNet-50 resnetv2_101 ResNet-V2-101
convmixer_1536_20 ConvMixer-1536/20 halonet50ts HaloNet-50 resnetv2_50 ResNet-V2-50
convnext_tiny ConvNeXt-T haloregnetz_b HaloRegNetZ resnetv2_50d_evos ResNet-V2-50-EVOS
crossvit_base_240 CrossViT-B hrnet_w64 HRNet-W32 resnext50_32x4d ResNeXt-50
cspdarknet53 CSPDarkNet-53 jx_nest_tiny NesT-T rexnet_150 ReXNet ($\times$1.5)
cspresnet50 CSPResNet-50 lambda_resnet50ts LambdaResNet-50 sebotnet33ts_256 SEBoTNet-33
cspresnext50 CSPResNeXt-50 lamhalobotnet50ts_256 LamHaloBoTNet-50 sehalonet33ts SEHaloNet-33
deit_base_patch16_224 DeiT-B mixnet_xl MixNet-XL seresnet50 SEResNet-50
deit_small_patch16_224 DeiT-S nf_regnet_b1 NF-RegNet-B1 seresnext50_32x4d SEResNeXt-50
dla102x2 DLA-X-102 nf_resnet50 NF-ResNet-50 swin_s3_tiny_224 S3 (Swin-T)
dpn107 DPN-107 nfnet_l0 NFNet-L0 swin_tiny_patch4_window7_224 Swin-T
eca_botnext26ts_256 ECA-BoTNeXt-26 pit_b_224 PiT-B tnt_s_patch16_224 TNT-S
eca_halonext26ts ECA-HaloNeXt-26 pit_s_224 PiT-S twins_pcpvt_base Twins-PCPVT-B
eca_nfnet_l0 ECA-NFNet-L0 poolformer_m48 PoolFormer-M48 twins_svt_small Twins-SVT-S
eca_resnet33ts ECA-ResNet-33 regnetx_320 RegNetX-320 visformer_small VisFormer-S
efficientnet_b2 EfficientNet-B2 regnety_032 RegNetY-32 vit_base_patch32_224 ViT-B
ese_vovnet39b eSE-VoVNet-39 resmlp_24_224 ResMLP-S24 vit_small_patch16_224 ViT-S
fbnetv3_g FBNetV3-G resmlp_big_24_224 ResMLP-B24 vit_small_r26_s32_224 R26+ViT-S
gcresnet50t GCResNet-50 resnest50d ResNeSt-50 wide_resnet50_2 Wide ResNet-50
gcresnext50ts GCResNeXt-50 resnet101 ResNet-101 xcit_medium_24_p16_224 XCiT-M24
gluon_resnet101_v1c ResNet-101-C resnet50 ResNet-50 xcit_tiny_12_p8_224 XCiT-T12
Table A.1: Lists of 69 models and their names based on their research paper and timm library.
Components Elements
Base architecture CNN, Transformer, MLP-Mixer, Hybrid (CNN + Transformer), NAS-Net
Stem layer 7$\times$7 conv with stride 2, 3$\times$3 conv with stride 2, 16$\times$16 conv with stride 16, …
Input resolution 224$\times$224, 256$\times$256, 240$\times$240, 299$\times$299
Normalization layer BN, GN, LN, LN + GN, LN + BN, Normalization-free, …
Using hierarchical structure Yes (e.g., CNNs, Swin [65]), No (e.g., ViT [28])
Activation functions ReLU, HardSwish, SiLU, GeLU, ReLU + GeLU, ReLU + SiLU or GeLU …
Using pooling at stem Yes, No
Using 2D self-attention Yes (e.g., [5], [87], [102]), No
Using channel-wise (CW) attention Yes (e.g., [46], [10], [106]), No
Using depth-wise convolution Yes, No
Using group convolution Yes, No
Type of pooling for final feature Classification (CLS) token, Global Average Pool (GAP)
Location of CW attentions At the end of each block, in the middle of each block, …
Table A.2: Overview of model elements. We categorize each architecture with 13 different architectural components. The full feature list of each architecture is in Tab. A.3 and Tab. A.4.
Table A.3: Description of features of 69 models. “s” in “Stem layer” indicates the stride of a layer in the stem, and the number before and after “s” are a kernel size and size of stride, respectively. For example, “3s2/3/3” means that the stem is composed of the first layer having 3 $\times$ 3 kernel with stride 2, the second layer having 3 $\times$ 3 kernel with stride 1, and the last layer having 3 $\times$ 3 with stride 1.
Model name Base architecture Hierarchical structure Stem layer Input resolution Normalization Activation
botnet26t_256 CNN Yes 3s2/3/3 256 $\times$ 256 BN ReLU
convmixer_1536_20 CNN Yes 7s7 224 $\times$ 224 BN GeLU
convnext_tiny CNN Yes 4s4 224 $\times$ 224 LN GeLU
cspdarknet53 CNN Yes 3s1 256 $\times$ 256 BN Leaky ReLU
cspresnet50 CNN Yes 7s2 256 $\times$ 256 BN Leaky ReLU
cspresnext50 CNN Yes 7s2 256 $\times$ 256 BN Leaky ReLU
dla102x2 CNN Yes 7s1 224 $\times$ 224 BN ReLU
dpn107 CNN Yes 7s2 224 $\times$ 224 BN ReLU
eca_botnext26ts_256 CNN Yes 3s2/3/3 256 $\times$ 256 BN SiLU
eca_halonext26ts CNN Yes 3s2/3/3 256 $\times$ 256 BN SiLU
eca_nfnet_l0 CNN Yes 3s2/3/3/3s2 224 $\times$ 224 Norm-free SiLU
eca_resnet33ts CNN Yes 3s2/3/3s2 256 $\times$ 256 BN SiLU
ese_vovnet39b CNN Yes 3s2/3/3s2 224 $\times$ 224 BN ReLU
gcresnet50t CNN Yes 3s2/3/3s2 256 $\times$ 256 LN + BN ReLU
gcresnext50ts CNN Yes 3s2/3/3 256 $\times$ 256 LN + BN ReLU + SiLU
gluon_resnet101_v1c CNN Yes 3s2/3/3 224 $\times$ 224 BN ReLU
gluon_xception65 CNN Yes 3s2/3 299 $\times$ 299 BN ReLU
halo2botnet50ts_256 CNN Yes 3s2/3/3s2 256 $\times$ 256 BN SiLU
halonet50ts CNN Yes 3s2/3/3 256 $\times$ 256 BN SiLU
hrnet_w64 CNN Yes 3s2/3s2 224 $\times$ 224 BN ReLU
lambda_resnet50ts CNN Yes 3s2/3/3 256 $\times$ 256 BN SiLU
lamhalobotnet50ts_256 CNN Yes 3s2/3/3s2 256 $\times$ 256 BN SiLU
nf_resnet50 CNN Yes 7s2 256 $\times$ 256 Norm-free ReLU
nfnet_l0 CNN Yes 3s2/3/3/3s2 224 $\times$ 224 Norm-free ReLU + SiLU
poolformer_m48 CNN Yes 7s4 224 $\times$ 224 GN GeLU
resnest50d CNN Yes 3s2/3/3 224 $\times$ 224 BN ReLU
resnet101 CNN Yes 7s2 224 $\times$ 224 BN ReLU
resnet50 CNN Yes 7s2 224 $\times$ 224 BN ReLU
resnet50_gn CNN Yes 7s2 224 $\times$ 224 GN ReLU
resnetblur50 CNN Yes 7s2 224 $\times$ 224 BN ReLU
resnetv2_101 CNN Yes 7s2 224 $\times$ 224 BN ReLU
resnetv2_50 CNN Yes 7s2 224 $\times$ 224 BN ReLU
resnetv2_50d_evos CNN Yes 3s2/3/3 224 $\times$ 224 EvoNorm -
resnext50_32x4d CNN Yes 7s2 224 $\times$ 224 BN ReLU
sebotnet33ts_256 CNN Yes 3s2/3/3s2 256 $\times$ 256 BN ReLU + SiLU
sehalonet33ts CNN Yes 3s2/3/3s2 256 $\times$ 256 BN ReLU + SiLU
seresnet50 CNN Yes 7s2 224 $\times$ 224 BN ReLU
seresnext50_32x4d CNN Yes 7s2 224 $\times$ 224 BN ReLU
wide_resnet50_2 CNN Yes 7s2 224 $\times$ 224 BN ReLU
convit_base Transformer No 16s16 224 $\times$ 224 LN GeLU
crossvit_base_240 Transformer Yes 16s16 240 $\times$ 240 LN GeLU
deit_base_patch16_224 Transformer No 16s16 224 $\times$ 224 LN GeLU
deit_small_patch16_224 Transformer No 16s16 224 $\times$ 224 LN GeLU
jx_nest_tiny Transformer Yes 4s4 224 $\times$ 224 LN GeLU
pit_s_224 Transformer Yes 16s8 224 $\times$ 224 LN GeLU
swin_tiny_patch4_window7_224 Transformer Yes 4s4 224 $\times$ 224 LN GeLU
tnt_s_patch16_224 Transformer Yes 7s4 224 $\times$ 224 LN GeLU
vit_base_patch32_224 Transformer No 32s32 224 $\times$ 224 LN GeLU
vit_small_patch16_224 Transformer No 16s16 224 $\times$ 224 LN GeLU
gmlp_s16_224 MLP-Mixer Yes 16s16 224 $\times$ 224 LN GeLU
resmlp_24_224 MLP-Mixer No 16s16 224 $\times$ 224 Affine transform GeLU
resmlp_big_24_224 MLP-Mixer Yes 8s8 224 $\times$ 224 Affine transform GeLU
swin_s3_tiny_224 NAS (TFM) Yes 4s4 224 $\times$ 224 LN GeLU
efficientnet_b2 NAS (CNN) Yes 3s2 256 $\times$ 256 BN SiLU
fbnetv3_g NAS (CNN) Yes 3s2 240 $\times$ 240 BN HardSwish
haloregnetz_b NAS (CNN) Yes 3s2 224 $\times$ 224 BN ReLU + SiLU
mixnet_xl NAS (CNN) Yes 3s2 224 $\times$ 224 BN ReLU + SiLU
nf_regnet_b1 NAS (CNN) Yes 3s2 256 $\times$ 256 Norm-free ReLU + SiLU
regnetx_320 NAS (CNN) Yes 3s2 224 $\times$ 224 BN ReLU
regnety_032 NAS (CNN) Yes 3s2 224 $\times$ 224 BN ReLU
rexnet_150 NAS (CNN) Yes 3s2 224 $\times$ 224 BN ReLU + SiLU + ReLU6
coat_lite_small Hybrid Yes 4s4 224 $\times$ 224 LN GeLU
pit_b_224 Hybrid Yes 14s7 224 $\times$ 224 LN GeLU
twins_pcpvt_base Hybrid Yes 4s4 224 $\times$ 224 LN GeLU
twins_svt_small Hybrid Yes 4s4 224 $\times$ 224 LN GeLU
visformer_small Hybrid Yes 7s2 224 $\times$ 224 BN GeLU + ReLU
vit_small_r26_s32_224 Hybrid No 7s2 224 $\times$ 224 LN + GN GeLU + ReLU
xcit_medium_24_p16_224 Hybrid No 3s2/3s2/3s2/3s2 224 $\times$ 224 LN + BN GeLU
xcit_tiny_12_p8_224 Hybrid No 3s2/3s2/3s2 224 $\times$ 224 LN + BN GeLU
Table A.4: Description of features of 69 models. “Pooling (stem)” and “Pooling (final)” denote “Pooling at the stem” and “Pooling for final feature”, respectively. “SA”, “CW”, and “DW” means “Self-attention”, “Channel-wise”, and “Depth-wise”, respectively.
Model name Pooling (stem) Pooling (final) 2D SA CW attention Location of CW attention DW conv Group conv
botnet26t_256 Yes GAP Yes (BoT) No No No
convmixer_1536_20 No GAP No No Yes No
convnext_tiny No GAP No No Yes No
cspdarknet53 No GAP No No No No
cspresnet50 Yes GAP No No No No
cspresnext50 Yes GAP No No No Yes
dla102x2 No GAP No No No Yes
dpn107 Yes GAP No No No Yes
eca_botnext26ts_256 Yes GAP Yes (BoT) Yes (ECA) Middle No Yes
eca_halonext26ts Yes GAP Yes (Halo) Yes (ECA) Middle No Yes
eca_nfnet_l0 No GAP No Yes (ECA) End No Yes
eca_resnet33ts No GAP No Yes (ECA) Middle No No
ese_vovnet39b No GAP No Yes (ESE) End No No
gcresnet50t No GAP No Yes (GCA) Middle Yes No
gcresnext50ts Yes GAP No Yes (GCA) Middle Yes Yes
gluon_resnet101_v1c Yes GAP No No No No
gluon_xception65 No GAP No No Yes No
halo2botnet50ts_256 No GAP Yes (Halo, BoT) No No No
halonet50ts Yes GAP Yes (Halo) No No No
hrnet_w64 No GAP No No No No
lambda_resnet50ts Yes GAP Yes (Lambda) No No No
lamhalobotnet50ts_256 No GAP Yes (Lambda, Halo, BoT) No No No
nf_resnet50 Yes GAP No No No No
nfnet_l0 No GAP No Yes (SE) End No Yes
poolformer_m48 No GAP No No No No
resnest50d Yes GAP No Yes No Yes
resnet101 Yes GAP No No No No
resnet50 Yes GAP No No No No
resnet50_gn Yes GAP No No No No
resnetblur50 Yes GAP No No No No
resnetv2_101 Yes GAP No No No No
resnetv2_50 Yes GAP No No No No
resnetv2_50d_evos Yes GAP No No No No
resnext50_32x4d Yes GAP No No No Yes
sebotnet33ts_256 No GAP Yes (BoT) Yes (SE) Middle No No
sehalonet33ts No GAP Yes (Halo) Yes (SE) Middle No No
seresnet50 Yes GAP No Yes (SE) End No No
seresnext50_32x4d Yes GAP No Yes (SE) End No Yes
wide_resnet50_2 Yes GAP No No No No
convit_base No CLS token No No No No
crossvit_base_240 No CLS token No No No No
deit_base_patch16_224 No CLS token No No No No
deit_small_patch16_224 No CLS token No No No No
jx_nest_tiny No GAP No No No No
pit_s_224 No CLS token No No Yes No
swin_tiny_patch4_window7_224 No GAP No No No No
tnt_s_patch16_224 No CLS token No No No No
vit_base_patch32_224 No CLS token No No No No
vit_small_patch16_224 No CLS token No No No No
gmlp_s16_224 No GAP No No No No
resmlp_24_224 No GAP No No No No
resmlp_big_24_224 No GAP No No No No
swin_s3_tiny_224 No GAP No No No No
efficientnet_b2 No GAP No Yes (SE) Middle Yes No
fbnetv3_g No GAP No Yes (SE) Middle Yes No
haloregnetz_b No GAP Yes (Halo) Yes (SE) Middle No Yes
mixnet_xl No GAP No Yes (SE) Middle Yes No
nf_regnet_b1 No GAP No Yes (SE) Middle Yes Yes
regnetx_320 No GAP No No No Yes
regnety_032 No GAP No Yes (SE) Middle No Yes
rexnet_150 No GAP No Yes (SE) Middle Yes No
coat_lite_small No CLS token No No Yes No
pit_b_224 No CLS token No No Yes No
twins_pcpvt_base No GAP No No Yes No
twins_svt_small No GAP No No Yes No
visformer_small No GAP No No No Yes
vit_small_r26_s32_224 Yes CLS token No No No No
xcit_medium_24_p16_224 No CLS token No No Yes No
xcit_tiny_12_p8_224 No CLS token No No Yes No
Appendix B Details of feature importance analysis
We fit a gradient boosting regressor [30] based on Scikit-learn [75] and report the permutation importance of each architectural component in Fig. 3 of the main paper. The number of boosting stages, maximum depth, minimum number of samples, and learning rate are set to 500, 12, 4, and 0.02, respectively. Permutation importance is computed by permuting a feature 10 times.
Appendix C Spectral clustering details
Figure C.1: Spectral features of 69 architectures. The $K$-th largest eigenvectors of the Laplacian matrix of the pairwise similarity graph of 69 architectures are shown ($K=10$ in this figure). Rows and columns are sorted by the clustering index in Tab. 2. We denote the model name in timm for each row and column.
We use the normalized Laplacian matrix to compute the Laplacian matrix. We also run K-means clustering 100 times and choose the clustering result with the best final objective function to reduce the randomness by the K-means clustering algorithm.
We visualize the pairwise distances of the spectral features (i.e., $K$-largest eigenvectors of $L$) of 69 architectures with their model names in Fig. C.1. This figure is an extension of Fig. 4(a) of the main paper, now including the model names. Note that rows and columns of Fig. C.1 are sorted by the clustering results. Fig. C.1 shows block diagonal patterns, i.e., in-cluster similarities are large while between-cluster similarities are small.
Appendix D Pairwise similarities
Figure D.1: Pairwise similarity among 69 models. Rows and columns are sorted by the clustering index in Tab. 2. ($n$, $n$) component of pairwise similarity is close to 4.6 ($\log 100$) because the attack success rate is almost 100% when a model used for generating adversarial perturbation and attacked model are the same.
Fig. D.1 indicates the pairwise similarity among 69 models. We can observe a weak block pattern around clusters, as also revealed in Fig. C.1.
Appendix E Training setting details for Sec. Training strategy and similarity score
We train 21 ResNet-50 models and 16 ViT-S from scratch individually by initializing each network with different random seeds. We further train 28 ResNet-50 models by randomly choosing learning rate ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base learning rate is 0.1), weight decay ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base weight decay is 1e-4), and learning rate scheduler (step decay or cosine decay). Similarly, we train 9 ViT-S models by randomly choosing learning rate ($\times$0.2, $\times$0.4, and $\times$1 where the base learning rate is 5e-4) and weight decay ($\times$0.2, $\times$0.4, and $\times$1 where the base weight decay is 0.05). Note that the DeiT training is unstable when we use a larger learning rate or weight decay than the base values. Finally, we collect 22 ResNet-50 models with different training regimes: 1 model with standard training by PyTorch [74]; 4 models trained by GluonCV [37]111gluon_resnet50_v1b, gluon_resnet50_v1c, gluon_resnet50_v1d, and gluon_resnet50_v1s from timm library.; a semi-supervised model and semi-weakly supervised model on billion-scale unlabeled images by Yalniz et al. [111]222ssl_resnet50 and swsl_resnet50 from timm library.; 5 models trained by different augmentation methods (Cutout [26], Mixup [117], manifold Mixup [103], CutMix [114], and feature CutMix [^3] We use the official weights provided by https://github.com/clovaai/CutMix-PyTorch.; 10 optimized ResNet models by [108] [^4] We use the official weights provided by https://github.com/rwightman/pytorch-image-models/releases/tag/v0.1-rsb-weights. We also collect 7 ViT-S models with different training regimes, including the original ViT training setup [28]555deit_small_patch16_224, vit_small_patch16_224, deit_small_distilled_patch16_224, and deit3_small_patch16_224 from timm library., a stronger data augmentation setup in the Deit paper [96]footnotemark: , the training setup with distillation [96]footnotemark: , an improved DeiT training setup [98]footnotemark: , and self-supervised training fashions by MoCo v3 [13] [^6] We train the ViT-S model by following https://github.com/facebookresearch/moco-v3, MAE [42] [^7] We train the ViT-S model by following https://github.com/facebookresearch/mae and BYOL [35] [^8] We train the ViT-S model by following https://github.com/lucidrains/byol-pytorch. We do not use adversarially-trained networks because the adversarial training usually drops the standard accuracy significantly [100].
Appendix F Ensemble performance
Fig. F.1 shows the ensemble performance by ImageNet top-1 error (lower is better). It has the same tendency as the error reduction rate reported in Fig. 6 of the main paper.
Figure F.1: Model diversity by architecture difference and ImageNet top-1 error ensemble performance. We report ImageNet top-1 error by varying the number of ensembled models ($N$) and the diversity of the ensembled models. The diversity is controlled by choosing the models from $k$ different clusters (larger $k$ denotes more diversity).
Appendix G Knowledge distillation
G.1 Training setting details for Sec. Model diversity and knowledge distillation
We train ViT-Ti student models with 25 different teacher models using hard distillation strategy. We follow the distillation training setting of DeiT official repo [^9] https://github.com/facebookresearch/deit., only changing the teacher model. Note that we resize the input images to the input size the teacher model requires if the input sizes of student and teacher models differ. If a teacher model needs a different input resolution, such as $240\times 240$ and $256\times 256$, we resize the input image for distilling it. Because DeiT-Ti has low classification accuracy compared to teacher models, the similarity score is calculated between DeiT-S and 25 models. The 25 teacher models are as follows: BoTNet-26, CoaT-Lite Small, ConViT-B, ConvNeXt-B, CrossViT-B, CSPDarkNet-53, CSPResNeXt-50, DLA-X-102, DPN-107, EfficientNet-B2, FBNetV3-G, GC-ResNet-50, gMLP-S, HaloRegNetZ, MixNet-XL, NFNet-L0, PiT-S, RegNetY-032, ResMLP-24, ResNet-50, SEHaloNet33, Swin-T, TNT-S, VisFormer-S, and XCiT-T12.
G.2 Teacher accuracy and distillation performance
Figure G.1: Teacher network accuracy and distillation performance. The trend line and its 90% confidence interval are shown in the blue line and gray band. There is no significant correlation between teacher accuracy and distillation performance.
The similarity between teacher and student networks may not be the only factor contributing to distillation. For example, a stronger teacher can lead to better distillation performance [115]. In Fig. G.1, we observe that if the range of the teacher accuracy is not significantly large enough (e.g., between 79.5 and 82.5), then the correlation between teacher network accuracy and distillation performance is not significant; 0.049 Pearson correlation coefficient with 0.82 p-value. In this case, we can conclude that the teacher and student networks similarity contributes more to the distillation performance than the teacher performance.
Appendix H Limtations and discussions
Figure H.1: Similarity scores using two-side attack transferability vs. one-side attack transferability. The trend line and its 90% confidence interval are shown in the blue line and gray band. We show the relationship between similarity scores using two-side and one-side attack transferability.
Similarity scores for adversarially trained networks.
Because our method is based on an adversarial attack, it could be affected by an adversarial attack-aware learning method. We first note that our analysis is invariant to the adversarial training; as shown by Tsipras et al. [100], adversarial robustness and accuracy are in a trade-off, and there is no ImageNet model with accuracy aligned with our target models yet. Also, our model similarity is based on the difference in the change of the model predictions. As shown by Tsipras et al. [100] and Ilyas et al. [49], adversarial training will lead to a different decision boundary. Therefore, we think that although our score can be affected by adversarial training, our score is still valid for adversarially trained models because an adversarially trained model is actually dissimilar to other models in terms of the decision boundary and representation learning.
An efficient approximation for a new model.
We can use our toolbox for designing a new model; we can measure the similarity between a novel network and existing $N$ architectures; a novel network can be assigned to clusters (Tab. 2) to understand how it works. However, it requires generating adversarial samples for all $N$+1 models (e.g., 70 in our case), which is computationally inefficient. Instead, we propose the approximation of Eq. (1) of the main paper by omitting to compute the accuracy of the novel network on the adversarial samples of the existing networks. It will break the symmetricity of our score, but we found that the approximated score and the original score have high similarity (0.82 Pearson correlation coefficient with almost 0 p-value) as shown in Fig. H.1).
As an example, we tested Gluon-ResNeXt-50 [37], which has the same architecture with ResNeXt-50 [110] in 69 models, and the distilled version of DeiT-S model [96]. Because the difference is not significant as we observed in Tab. 2 and Fig. 4 (b) of the main paper, we expect that Gluon-ResNeXt-50 is assigned to the same cluster with ResNeXt-50, and distilled DeiT-S is assigned to the same cluster with DeiT-S. We observe that, as we expected, each network is assigned to the desired cluster. Therefore, we suggest using our efficient approximation for analyzing a novel network with our analysis toolbox.
Other possible applications.
We can use our similarity to develop a novel neural architecture or training technique by measuring how the proposed method can learn different representations compared to existing methods. Our similarity also can be used for analyzing the hyper-parameter sensitivity. Instead of reporting the performance changes by different hyper-parameters, one can report the average similarity of the trained models with different hyper-parameters. If the average similarity is significantly low, then we can argue that the method is sensitive to hyper-parameter selection. We expect this scenario is especially useful for the task, which is sensitive to the hyper-parameter selection [16, 36]. Finally, our similarity can be used for the model selection criteria for selecting diverse models, e.g., constructing a new dataset with diverse pre-trained models for annotation candidates [59, 20].